A firm has sales of $500,000, a debt-to-equity ratio of one, and total assets of $1,000,000. If its profit margin is 5%, what is the firm’s return on equity?

a) 3.3%

b) 6.7 %

c) 5.0 %

d) 2.5 %

e) Further information is needed,
Answer #1

Profit margin=Net income/Sales

Net income=(500,000*5%)=$25000

debt to Equity ratio=debt/Equity

Hence debt=Equity

Total assets=Total liabilities+Total equity

1,000,000=Equity+Equity

Equity=1,000,000/2=$500,000

ROE=Net income/Equity

=25000/500,000

=5%